The Lumber Yard is considering adding a new product line that is expected to increase annual sales by $230,000 and cash expenses by $146,000. The initial investment will require $120,000 in fixed assets that will be depreciated using the 5-year MACRS. The company has a marginal tax rate of 29 percent. What is the project OCF in year 2? (Do not include the dollar sign ($). Round your answer to a whole dollar. (e.g., 4,132) MACRS 5-year property

Year Rate

1 20.00%

2 32.00%

3 19.20%

4 11.52%

5 11.52%

6 5.76%
Answer #1
Computation of operating cash flow in year 2
i sales = 230,000
ii Cash expenses 146,000
iii Depreciation 38400
=120,000*32%
iv=i-ii-iii Profit before tax 45,600
v Tax rate @ 29% 13224
vi=iv-v Profit after tax 32,376
vii=vi+iii Operating cash flow 70,776
Therefore answer = 70,776